OpenStudy (anonymous):

4 2/7 converted to a decimal and rounded to the nearest thousandths ? please help need the awnser asap

OpenStudy (anonymous):

\[4\tfrac{2}{7}=\frac{4\times 7+2}{7}=\frac{30}{7}\] take a calculator and compute \(30\div 7\) for your decimal

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=30%2F7

OpenStudy (anonymous):

thank you so much